OpenStudy (anonymous):

A rifle is aimed horizontally at a target 55.0 m away. The bullet hits the target 3.00 cm below the aim point What was the bullet's flight time? What was the bullet's speed as it left the barrel?

OpenStudy (anonymous):

Initial vertical velocity, Vy = 0. So, time of flight = time taken by bullet to fall 3 cm vertically. Use S = 1/2*a*t^2 t = sqrt.( 2s/a) t= sqrt.(2*0.03/9.8) sec.